# How Far Is Video Generation from World Model: A Physical Law Perspective

### v1

@@ Abstract
OpenAI’s Sora highlights the potential of video generation for developing world models that adhere to fundamental physical laws. However, the ability of video generation models to discover such laws purely from visual data without human priors can be questioned. A world model learning the true law should give predictions robust to nuances and correctly extrapolate on unseen scenarios. In this work, we evaluate across three key scenarios: in-distribution, out-of-distribution, and combinatorial generalization. We developed a 2D simulation testbed for object movement and collisions to generate videos deterministically governed by one or more classical mechanics laws. This provides an unlimited supply of data for large-scale experimentation and enables quantitative evaluation of whether the generated videos adhere to physical laws. We trained diffusion-based video generation models to predict object movements based on initial frames. Our scaling experiments show perfect generalization within the distribution, measurable scaling behavior for combinatorial generalization, but failure in out-of-distribution scenarios. Further experiments reveal two key insights about the generalization mechanisms of these models: (1) the models fail to abstract general physical rules and instead exhibit “case-based” generalization behavior, i.e., mimicking the closest training example; (2) when generalizing to new cases, models are observed to prioritize different factors when referencing training data: color size velocity shape. Our study suggests that scaling alone is insufficient for video generation models to uncover fundamental physical laws, despite its role in Sora’s broader success.
@@ 1 Introduction
Foundation models (Bommasani et al., 2021) have emerged remarkable capabilities by scaling the model and data to an unprecedented scale (Brown, 2020; Kaplan et al., 2020). As an example, OpenAI’s Sora (Brooks et al., 2024) not only generates high-fidelity and surreal videos, but also has sparked a new surge of interest in studying world models (Yang et al., 2023).
“Scaling video generation models is a promising path towards building general purpose simulators of the physical world.” — Sora Report (Brooks et al., 2024)
@@
World simulators are receiving broad attention from robotics (Yang et al., 2023) and autonomous driving (Hu et al., 2023) for the ability to generate realistic data and accurate simulations. These models are required to comprehend fundamental physical laws to produce data that extends beyond the training corpus and to guarantee precise simulation. However, it remains an open question whether video generation can discover such rules merely by observing videos, as Sora does. We aim to provide a systematic study to understand the critical role and limitation of scaling in physical law discovery.
@@
It is challenging to determine whether a video model has learned a law instead of merely memorizing the data. Since the model’s internal knowledge is inaccessible, we can only infer the model’s understanding by examining its predictions on unseen scenarios, i.e., its generalization ability. We propose a categorization (Figure 1) for comprehensive evaluation based on the relationship between training and testing data in this paper. In-distribution (ID) generalization assumes that training and testing data are independent and identically distributed (i.i.d.). Out-of-distribution (OOD) generalization, on the other hand, refers to the model’s performance on testing data that come from a different distribution than the training data, particularly when latent parameters fall outside the range seen during training. Human-level physical reasoning can easily extrapolate OOD and predict physical processes without encountering the exact same scenario before. Additionally, we also examine a special OOD capacity called combinatorial generalization, which assesses whether a model can combine two distinct concepts in a novel way, a trait often considered essential for foundation models in advancing toward artificial general intelligence (AGI) (Du & Kaelbling, 2024).
@@
Moreover, real-world videos typically contain complex, non-rigid objects and motions, which present significant challenges for quantitative evaluation and even human validation. The rich textures and appearances in such videos can act as confounding factors, distracting the model from focusing on the underlying physics. To mitigate these issues, we specifically focus on classical mechanics and develop a 2D simulator with objects represented by simple geometric shapes. Each video depicts the motion or collision of these 2D objects, governed entirely by one or two fundamental physical laws, given the initial frames. This simulator allows us to generate large-scale datasets to support the scaling of video generation models. Additionally, we have developed a tool to infer internal states (e.g., the position and size) of each object in the generated video from pixels. This enables us to establish quantitative evaluation metrics for physical law discovery.
@@
We begin by investigating how scaling video generation models affects ID and OOD generalization. We select three fundamental physical laws for simulation: uniform linear motion of a ball, perfectly elastic collision between two balls, and parabolic motion of a ball. We scale the dataset from 30K to 3 million examples and increase the video diffusion model’s parameters from 22M to 310M. Consistently, we observe that the model achieves near-perfect ID generalization across all tasks. However, the OOD generalization error does not improve with increased data and model size, revealing the limitations of scaling video generation models in handling OOD data. For combinatorial generalization, we design an environment that involves multiple objects undergoing free fall and collisions to study their interactions. Every time, four objects from eight are selected to create a video. In total, 70 combinations () are possible. We use 60 of them for training and 10 for testing. We train models by varying the number of training data from 600K to 6M. We manually evaluate the generated test samples by labeling them as “abnormal” if the video looks physically implausible. The results demonstrate that scaling the data substantially reduces the percentage of abnormal cases, from to . This suggests that scaling is critical for improving combinatorial generalization.
@@
Our empirical analysis reveals two intriguing properties of the generalization mechanism in video generation models. First, these models can be easily biased by "deceptive" examples from the training set, leading them to generalize in a "case-based" manner under certain conditions. This phenomenon, also observed in large language models (Hu et al., 2024), describes a model’s tendency to reference similar training cases when solving new tasks. For instance, consider a video model trained on data of a high-speed ball moving in uniform linear motion. If data augmentation is performed by horizontally flipping the videos, thereby introducing reverse-direction motion, the model may generate a scenario where a low-speed ball reverses direction after the initial frames, even though this behavior is not physically correct. Second, we explore how different data attributes compete during the generalization process. For example, if the training data for uniform motion consists of red balls and blue squares, the model may transform a red square into a ball immediately after the conditioning frames. This behavior suggests that the model prioritizes color over shape. Our pairwise analysis reveals the following prioritization hierarchy: color > size > velocity > shape. This ranking could explain why current video generation models often struggle with maintaining object consistency.
We hope these findings provide valuable insights for future research in video generation and the development of world models.
2 Discovering Physics Laws with Video Generation
@@ 2.1 Problem Definition
Consider a physical procedure which involves several latent variables , each standing for a certain physical parameter such as velocity or position. By classical mechanics, these latent variables will evolve by differential equation . In discrete version, if time gap between two consecutive frames is , then we have . Denote rendering function as which render the state of the world into an image of shape with RGB channels. Consider a video consisting of frames that follows the classical mechanics dynamics. The physical coherence requires that there exists a series of latent variables which satisfy following requirement: 1) 2) We train a video generation model parametried by , where characterizes its understanding of video frames. We can predict the subsequent frames by sampling from based on initial frames’ condition. The variable usually takes the value of 1 or 3 depends on tasks. Therefore, physical-coherence loss can be simply defined as . It measures how likely the predicted value will cater to the real world development. The model must understand the underlying physical process to accurately forecast subsequent frames, which we can quantatively evaluate whether video generation model correctly discover and simulate the physical laws.
@@ 2.2 Video Generation Model
Following Sora (Brooks et al., 2024), we adopt the Variational Auto-Encoder (VAE) and DiT architectures for video generation. The VAE compresses videos into latent representations both spatially and temporally, while the DiT models the denoising process. This approach demonstrates strong scalability and achieves promising results in generating high-quality videos.
VAE Model. We employ a (2+1)D-VAE to project videos into a latent space. Starting with the SD1.5-VAE structure, we extend it into a spatiotemporal autoencoder using 3D blocks (Yu et al., 2023b). All parameters of the (2+1)D-VAE are pretrained on high-quality image and video data to maintain strong appearance modeling while enabling motion modeling. More details are provided in Section A.3.1. In this paper, we fix the pretrained VAE encoder and use it as a video compressor. Results in Section A.3.2 confirm the VAE’s ability to accurately encode and decode the physical event videos. This allows us to focus solely on training the diffusion model to learn the physical laws.
Diffusion model. Given the compressed latent representation from the VAE model, we flatten it into a sequence of spacetime patches, as transformer tokens. Notably, self-attention is applied to the entire spatio-temporal sequence of video tokens, without distinguishing between spatial and temporal dimensions. For positional embedding, a 3D variant of RoPE (Su et al., 2024) is adopted. As stated in Sec.2.1, our video model is conditioned on the first frames. The -frame video is zero-padded to the same length as the full physical video. We also introduce a binary mask “video” by setting the value of the first frames to , indicating those frames are the condition inputs. The noise, condition and mask videos are concatenated along the channel dimension to form the final input to the model.
@@ 2.3 On the Verification of Learned Laws
Suppose we have a video generation model learned based on the above formulation. How do we determine if the underlying physical law has been discovered? A well-established law describes the behavior of the natural world, e.g., how objects move and interact. Therefore, a video model incorporating true physical laws should be able to withstand experimental verification, producing reasonable predictions under any circumstances, which demonstrates the model’s generalization ability. To comprehensively evaluate this, we consider the following categorization of generalization (see Figure 1) within the scope of this paper: 1) In-distribution (ID) generalization describes the setting where training data and testing data are from the same distribution. In our case, both training and testing data follow the same law and are located in the same domain. 2) A human who has learned a physical law can easily extrapolate to scenarios that have never been observed before. This ability is referred to as out-of-distribution (OOD) generalization. Although it sounds challenging, this evaluation is necessary as it indicates whether a model can learn principled rules from data. 3) Moreover, there is a situation between ID and OOD, which has more practical value. We call this combinatorial generalization, representing scenarios where every "concept" or object has been observed during training, but not their every combination. It examines a model’s ability to effectively combine relevant information from past experiences in novel ways. A similar concept has been explored in LLMs (Riveland & Pouget, 2024), which demonstrated that models can excel at linguistic instructing tasks by recombining previously learned components, without task-specific experience.
@@ 3 In-Distribution and Out-of-Distribution Generalization
In this section, we study how in-distribution and out-of-distribution generalization is correlated with model or data scaling. We focus on deterministic tasks governed by basic kinematic equations, as they allow clear definitions of ID/OOD and straightforward quantitative error evaluation.
@@ 3.1 Fundamental Physical Scenarios
Specifically, we consider three physical scenarios illustrated in Fig. 2. 1) Uniform Linear Motion: A colored ball moves horizontally with a constant velocity. This is used to illustrate the law of Intertia. 2) Perfectly Elastic Collision: Two balls with different sizes and speeds move horizontally toward each other and collide. The underlying physical law is the conservation of energy and momentum. 3) Parabolic Motion: A ball with a initial horizontal velocity falls due to gravity. This represents Newton’s second law of motion. Each motion is determined by its initial frames.
Training data generation. We use Box2D to simulate kinematic states for various scenarios and render them as videos, with each scenario having 2-4 degrees of freedom (DoF), such as the balls’ initial velocity and mass. An in-distribution range is defined for each DoF. We generate training datasets of 30K, 300K, and 3M videos by uniformly sampling a high-dimensional grid within these ranges. All balls have the same density, so their mass is inferred from their size. Gravitational acceleration is constant in parabolic motion for consistency. Initial ball positions are randomly initialized within the visible range. Further details are provided in Section A.4.
@@
Test data generation. We evaluate the trained model using both ID and OOD data. For ID evaluation, we sample from the same grid used during training, ensuring that no specific data point is part of the training set. OOD evaluation videos are generated with initial radius and velocity values outside the training range. There are various types of OOD setting, e.g. velocity/radius-only or both OOD. Details are provided in Section A.4.
Models. For each scenario, we train models of varying sizes from scratch, as shown in Table 1. This ensures that the outcomes are not influenced by uncontrollable pretrain data. The first three frames are provided as conditioning, which is sufficient to infer the velocity of the balls and predict the subsequent frames. Diffusion model is trained for 100K steps using 32 Nvidia A100 GPUs with a batch size of 256, which was sufficient for convergence, as a model trained for 300K steps achieves a similar performance. We keep the pretrained VAE fixed. Each video consists of 32 frames with a resolution of 128x128. We also experimented with a 256x256 resolution, which yielded a similar generalization error but significantly slowed down the training process.
@@
Evaluation metrics. We observed that the learned models are able to generate balls with consistent shapes. To obtain the center positions of the -th ball in the generated videos, , we use a heuristic algorithm based on the mean of colored pixels, distinguishing the balls by color. To ensure the correctness of , we exclude frames with part of a ball out of view, yielding valid frames . For collision scenarios, only the frames after the collision are considered. We then compute the velocity of each ball, , at each moment by differentiating their positions. The error for a video is defined as: where is the computed velocity at time , is the ground-truth velocity in the simulator, is the number of balls, and is the number of valid frames.
@@
Baseline. We calculate the error between the ground truth velocity and the parsed values from the ground truth video, referred to as Groundtruth (GT). This represents the system error—caused by parsing video into velocity—and defines the minimum error a model can achieve.
3.2 Main Result of Scaling data and model
@@
For in-distribution (ID) generalization in Figure 3, increasing the model size (DiT-S to DiT-L) or the data amount (30K to 3M) consistently decreases the velocity error across all three tasks, strongly evidencing the importance of scaling for ID generalization. Take the uniform motion task as an example: the DiT-S model has a velocity error of with 30K data, while DiT-L achieves an error of with 3M data, very close to the error of obtained with ground truth video.
@@
However, the results differ significantly for out-of-distribution (OOD) predictions. First, OOD velocity errors are an order of magnitude higher than ID errors in all settings. For example, the OOD error for the DiT-L model on uniform motion with 3M data is , while the ID error is just . Second, scaling up the training data and model size has little or negative impact on reducing this prediction error. The variation in velocity error is higly random as data or model size changes, e.g., the error for DiT-B on uniform motion is , and , with data amounts of 30K, 300K and 3M. We also trained DiT-XL on the uniform motion 3M dataset but observed no improvement in OOD generalization. As a result, we did not pursue training of DiT-XL on other scenarios or datasets constrained by resources. These findings suggest the inability of scaling to perform reasoning in OOD scenarios. The sharp difference between ID and OOD settings further motivates us to study the generalization mechanism of video generation in Section 5.2.

### v2

@@ Abstract
Scaling video generation models is believed to be promising in building world models that adhere to fundamental physical laws. However, whether these models can discover physical laws purely from vision can be questioned. A world model learning the true law should give predictions robust to nuances and correctly extrapolate on unseen scenarios. In this work, we evaluate across three key scenarios: in-distribution, out-of-distribution, and combinatorial generalization. We developed a 2D simulation testbed for object movement and collisions to generate videos deterministically governed by one or more classical mechanics laws. We focus on the scaling behavior of training diffusion-based video generation models to predict object movements based on initial frames. Our scaling experiments show perfect generalization within the distribution, measurable scaling behavior for combinatorial generalization, but failure in out-of-distribution scenarios. Further experiments reveal two key insights about the generalization mechanisms of these models: (1) the models fail to abstract general physical rules and instead exhibit “case-based” generalization behavior, i.e., mimicking the closest training example; (2) when generalizing to new cases, models are observed to prioritize different factors when referencing training data: color size velocity shape. Our study suggests that scaling alone is insufficient for video generation models to uncover fundamental physical laws.
booktabs
∗ Equal Contribution (in alphabetical order)
@@ 1 Introduction
Foundation models (Bommasani et al., 2021) have emerged remarkable capabilities by scaling the model and data to an unprecedented scale (Brown, 2020; Kaplan et al., 2020). As an example, video generation (Brooks et al., 2024; Yang et al., 2024) not only generates high-fidelity and surreal videos, but has also sparked a new surge of interest in the construction of world models (Yang et al., 2023). This topic is receiving great attention from robotics (Yang et al., 2023; Yue et al., 2024) and autonomous driving (Hu et al., 2023) due to the ability to generate realistic data and accurate simulations. These models are required to comprehend fundamental physical laws to produce data that extend beyond the training corpus and to guarantee precise simulation. However, it remains an open question whether video generation can discover such rules merely by observing videos. We aim to provide a systematic study to answer this question by specifically focusing on scaling up the data and models.
@@
It is challenging to determine whether a video model has learned a law instead of simply memorizing the data. Since the internal knowledge of the model is inaccessible, we can only infer the understanding of the model by examining its predictions on unseen scenarios, i.e., its generalization ability. In this paper, we propose a categorization (Figure 1) for a comprehensive evaluation based on the relationship between training and test data. In-distribution (ID) generalization assumes that training and testing data are independent and identically distributed (i.i.d.). Out-of-distribution (OOD) generalization, on the other hand, refers to the model’s performance on testing data that come from a different distribution than the training data, particularly when latent parameters fall outside the range seen during training. Human-level physical reasoning can easily extrapolate OOD and predict physical processes without having encountered the exact same scenario before. Additionally, we also examine a special OOD capacity called combinatorial generalization, which assesses whether a model can combine two distinct concepts in a novel way, a trait often considered essential for foundation models in the advancement toward artificial general intelligence (AGI) (Du & Kaelbling, 2024).
@@
Moreover, real-world videos typically contain complex, non-rigid objects and motions, which present significant challenges for quantitative evaluation and even human validation. The rich textures and appearances of such videos can act as confounding factors, distracting the model from focusing on the underlying physics. To mitigate these issues, we specifically focus on classical mechanics and develop a 2D simulator with objects represented by simple geometric shapes. Each video depicts the motion or collision of these 2D objects, governed entirely by one or two fundamental physical laws given the initial frames. This simulator allows us to generate large-scale datasets to support the scaling of video generation models. Additionally, we have developed a tool to infer internal states (e.g., the position and size) of each object in the generated video from pixels. This enables us to establish quantitative evaluation metrics for physical law discovery.
@@
We begin by investigating how scaling video generation models affects ID and OOD generalization. We select three fundamental physical laws for simulation: uniform linear motion of a ball, perfectly elastic collision between two balls, and parabolic motion of a ball. We scale the dataset from 30K to 3 million examples and increase the video diffusion model’s parameters from 22M to 310M. Consistently, we observe that the model achieves a near-perfect ID generalization across all tasks. However, the OOD generalization error does not improve with increased data and model size, revealing the limitations of scaling video generation models in handling OOD data. For combinatorial generalization, we design an environment that involves multiple objects undergoing free fall and collisions to study their interactions. Each time, four of the eight objects are selected to create a video. In total, 70 combinations () are possible. We use 60 of them for training and 10 for testing. We train models by varying the number of training data from 600K to 6M. We manually evaluate the generated test samples by labeling them as “abnormal” if the video looks physically implausible. The results demonstrate that scaling the data substantially reduces the percentage of abnormal cases from to . This suggests that scaling is critical for improving combinatorial generalization.
@@
Our empirical analysis reveals two intriguing properties of the generalization mechanism in video generation models. First, these models can be easily biased by ”deceptive” examples from the training set, leading them to generalize in a ”case-based” manner under certain conditions. This phenomenon, also observed in large language models (Hu et al., 2024), describes a model’s tendency to reference similar training cases when solving new tasks. For instance, consider a video model trained on data of a high-speed ball moving in uniform linear motion. If data augmentation is performed by horizontally flipping the videos, thereby introducing reverse-direction motion, the model may generate a scenario where a low-speed ball reverses direction after the initial frames, even though this behavior is not physically correct. Second, we explore how different data attributes compete during the generalization process. For example, if the training data for uniform motion consists of red balls and blue squares, the model may transform a red square into a ball immediately after the conditioning frames. This behavior suggests that the model prioritizes color over shape. Our pairwise analysis reveals the following prioritization hierarchy: color > size > velocity > shape. This ranking could explain why current video generation models often struggle with maintaining object consistency. We hope these findings provide valuable insights for future research in video generation and world models.
2 Video Generation for Phyiscal Law Discovery
@@ 2.1 Problem Definition
Consider a physical procedure which involves several latent variables , each representing a certain physical parameter such as velocity or position. By classical mechanics, these latent variables will evolve by a differential equation . In the discrete version, if the time gap between two consecutive frames is , then we have . Denote the rendering function as which renders the state of the world into an image of shape with RGB channels. Consider a video consisting of frames that follows the classical mechanics dynamics. The physical coherence requires that there exists a series of latent variables which satisfy the following requirement: 1) 2) We train a video generation model parametried by , where characterizes its understanding of video frames. We can predict the subsequent frames by sampling from based on initial frames’ condition. The variable usually takes the value of 1 or 3 depending on the tasks. Therefore, physical coherence loss can be simply defined as . It measures how likely the predicted value will cater to real-world development. The model must understand the underlying physical process to accurately forecast subsequent frames, which we can quantatively evaluate whether video generation models correctly discover and simulate the physical laws.
@@ 2.2 Video Generation Model
In this paper, we focus exclusively on video generation models, leaving discussions on other modeling methods for future work. Following Sora (Brooks et al., 2024), we adopt the Variational Auto-Encoder (VAE) and DiT architectures for video generation. The VAE compresses videos into latent representations both spatially and temporally, while the DiT models the denoising process. This approach shows strong scalability and achieves promising results in generating high-quality videos.
VAE Model. We employ a (2+1)D-VAE to project videos into a latent space. Starting with the SD1.5-VAE structure, we extend it into a spatiotemporal autoencoder using 3D blocks (Yu et al., 2023b). All parameters of the (2+1)D-VAE are pretrained on high-quality image and video data to maintain strong appearance modeling while enabling motion modeling. More details are provided in Section B.2. In this paper, we fix the pretrained VAE encoder and use it as a video compressor. The results in Section B.3 confirm the VAE’s ability to accurately encode and decode the physical event videos. This allows us to focus solely on training the diffusion model to learn the physical laws.
Diffusion model. Given the compressed latent representation of the VAE model, we flatten it into a sequence of spacetime patches, as transformer tokens. Notably, self-attention is applied to the entire spatio-temporal sequence of video tokens, without distinguishing between spatial and temporal dimensions. For positional embedding, a 3D variant of RoPE (Su et al., 2024) is adopted. As stated in Sec.2.1, our video model is conditioned on the first frames. The -frame video is zero-padded to the same length as the full physical video. We also introduce a binary mask “video” by setting the value of the first frames to , indicating that these frames are condition inputs. The noise, condition and mask videos are concatenated along the channel dimension to form the final input to the model.
@@ 2.3 On the Verification of Learned Laws
Suppose we have a video generation model learned as described above. How do we determine whether the underlying physical law has been discovered? A well-established law describes the behavior of the natural world, e.g., how objects move and interact. Therefore, a video model incorporating true physical laws should be able to withstand experimental verification, producing reasonable predictions under any circumstances, which demonstrates the model’s generalization ability. To comprehensively evaluate this, we consider the following generalization categorization (see Figure 1) within the scope of this paper: 1) In-distribution (ID) generalization describes the setting where training data and testing data are from the same distribution. In our case, both training and testing data follow the same law and are located in the same domain. 2) A human who has learned a physical law can easily extrapolate to scenarios that have never been observed before. This ability is referred to as out-of-distribution (OOD) generalization. Although it sounds challenging, this evaluation is necessary as it indicates whether a model can learn principled rules from data. 3) Moreover, there is a situation between ID and OOD, which has more practical value. We call this combinatorial generalization, representing scenarios where every “concept” or object has been observed during training, but not all of their combinations. It tests a model’s ability to combine relevant information from past experiences in novel ways. A similar concept has been explored in LLMs (Riveland & Pouget, 2024), which showed that models can excel at linguistic-instructing tasks by recombining previously learned components, without task-specific experience.
@@ 3 In-Distribution and Out-of-Distribution Generalization
In this section, we study the correlation between ID/OOD generalization and model or data scaling, focusing on video generation models. We use deterministic tasks governed by basic kinematic equations, as they allow clear definitions of ID/OOD and straightforward quantitative error evaluation.
@@ 3.1 Fundamental Physical Scenarios
Specifically, we consider three physical scenarios. 1) Uniform Linear Motion: A colored ball moves horizontally with a constant velocity. This is used to illustrate the law of Intertia. 2) Perfectly Elastic Collision: Two balls of different sizes and speeds move horizontally toward each other and collide. The underlying physical law is the conservation of energy and momentum. 3) Parabolic Motion: A ball with an initial horizontal velocity falls due to gravity. This represents Newton’s second law of motion. Each motion is determined by its initial frames. See Figure 2 for visualizations.
Training data generation. We use Box2D to simulate kinematic states for various scenarios and render them as videos, with each scenario having 2-4 degrees of freedom (DoF), such as the balls’ initial velocity and mass. An in-distribution range is defined for each DoF. We generate training datasets of 30K, 300K, and 3M videos by uniformly sampling a high-dimensional grid within these ranges. All balls have the same density, so their mass is inferred from their size. Gravitational acceleration is constant in parabolic motion for consistency. The initial ball positions are initialized randomly within the visible range. More details are provided in Section C.1.
Test data generation. We evaluate the trained model using both ID and OOD data. For ID evaluation, we sample from the same grid used during training, ensuring that no specific data point is part of the training set. OOD evaluation videos are generated with initial radius and velocity values outside the training range. There are various types of OOD setting, e.g. velocity/radius-only or both OOD.
Models. For each scenario, we train models of varying sizes from scratch, as shown in Table 1. This ensures that the results are not influenced by uncontrollable pretraining data. The first three frames are provided as conditioning, which is sufficient to infer the velocity of the balls and predict the subsequent frames. The diffusion model is trained for 100K steps using 32 Nvidia A100 GPUs with a batch size of 256, which was sufficient for convergence, since a model trained for 300K steps achieves similar performance. We keep the pre-trained VAE fixed. Each video consists of 32 frames with a resolution of 128x128. We also experimented with a 256x256 resolution, which yielded a similar generalization error but significantly slowed down the training process.
@@
Evaluation metrics. We observed that the learned models are able to generate balls with consistent shapes. To obtain the center positions of the -th ball in the generated videos, , we use a heuristic algorithm based on the mean of colored pixels, distinguishing the balls by color. To ensure the correctness of , we exclude frames with part of a ball out of view, yielding valid frames . For collision scenarios, only frames after the collision are considered. We then compute the velocity of each ball, , at each moment by differentiating their positions. The error for a video is defined as: where is the calculated velocity at time , is the ground-truth velocity in the simulator, is the number of balls and is the number of valid frames.
@@
Baseline. We calculate the error between the ground truth velocity and the parsed values of the ground truth video, referred to as Groundtruth (GT). This represents the system error—caused by parsing video into velocity—and defines the minimum error a model can achieve.
3.2 Perfect ID and Failed OOD Generalization
@@
For in-distribution (ID) generalization in Figure 3, increasing the model size (DiT-S to DiT-L) or the amount of data (30K to 3M) consistently decreases the velocity error across the three tasks, strongly evidencing the importance of scaling for ID generalization. Take the uniform motion task as an example: the DiT-S model has a velocity error of with 30K data, while DiT-L achieves an error of with 3M data, very close to the error of obtained with ground truth video.
@@
However, the results differ significantly for out-of-distribution (OOD) predictions. First, OOD velocity errors are an order of magnitude higher than ID errors in all settings. For example, the OOD error for the DiT-L model in uniform motion with 3M data is , while the ID error is just . Second, scaling up the training data and model size has little or negative impact on reducing this prediction error. The variation in velocity error is highly random as data or model size changes, e.g., the error for DiT-B on uniform motion is , and , with data amounts of 30K, 300K and 3M. We also trained DiT-XL on the uniform motion 3M dataset but observed no improvement in OOD generalization. As a result, we did not carry out DiT-XL training in other scenarios. These findings suggest the inability of scaling to perform reasonably in OOD scenarios. The sharp difference between the ID and OOD settings motivates us to study the generalization mechanism of video generation in Section 5.2.